OpenStudy (anonymous):

A sequence has its first term equal to 8 and each term of the sequence is obtained by adding 6 to the previous term. If f( n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence?

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

f(1) = 6 and f(n ) = f(n - 1) + 8; n > 1 f(1) = 8 and f(n ) = f(n - 1) + 6; n > 1 f(1) = 8 and f(n ) = f(n - 1) + 6n; n > 1 f(1) = 6 and f(n ) = f(n - 1) + 8n; n > 1

OpenStudy (anonymous):

Choice (2) is correct. Since the first term is 8, so f(1) = 8, so you should have immediately eliminated choices A and D.

OpenStudy (anonymous):

Can you please help me?: http://openstudy.com/updates/526082dae4b0c78c62eade37